3:10PM 4/18/2017 Update: Rolling reboots on Owens have started to address GPFS errors occured late Friday. 

3:10PM 4/18/2017 Update: Rolling reboots on Owens have started to address this GPFS issue. 

2/26 Downtime Difficulties

Category: 
Resolution: 
Resolved

All systems should be functioning normally. Please report any remaining issues to OSC Help.

----
A number of systems are still experiencing problems after yesterday's downtime. Currently, the following systems are still offline:

  • Oakley (returned to service)

  • ARMSTRONG (returned to service)

  • license server (returned to service)

  • proj11 (returned to service)

  • proj12 (returned to service)

  • proj13 (returned to service)

  • proj14 (returned to service)

Glenn is operational. We will update as the situation changes.